how to write the equation of a line passing through (1,2) and parallel to
3x-y=-7

Found 2 solutions by rfer, Mr_Mathman:
Answer by rfer(16322)

y=mx+b
2=3*1+b
2=3+b
b=-1
-------------
y=3x-1

Answer by Mr_Mathman(5)

First solve the eqn for y
3x-y=-7 subtract 3x from both sides
-y=-3x-7 multiply both sides by -1
y=3x+7
Replace 7 with b and replace (x,y) with (1,2). So,
2=3(1)+b=3+b subtract 3 from both sides
b=-1
Remember that y=mx+b and that parallel lines have the same slope. So in the parallel line m=3 and b=-1. This gives:
y=3x-1